KpopAsianDramaFanYT:

The revenue each season from tickets at the theme part is represented by t(x) = 3x. The cost to pay the employees each season is represented by r(x) = (1.25)^x. Examine the graph of the combined function for total profit and estimate the profit after five seasons. Answers: A) 2 B) 5 C) 12 D) 18 The graph will be in the Reply section!

KpopAsianDramaFanYT:

1 attachment
KpopAsianDramaFanYT:

I Really need help

KpopAsianDramaFanYT:

DM me 2 help me with this Question

VIBE:

t(x)=3x is the Revenue function. r(x)=1.25x represents the cost function. As we know the benefit formula, we must determine the overall profit. P(X)=t(x)-r(x) - P(x)=3 - x-1.25x Profit = Revenue - Cost After five seasons, we must calculate the profit: P(5)=$239.94 p(5)=35-1.255 As a result, the profit is $239.94 after five seasons.